During the first part of a trip, a canoeist travels 35 miles at a certain speed. The canoeist travels 14 miles on the second part or the trip at a speed 5 mph slower. The total time for the trip is 2 hours. What is the speed on each part of the trip?
(not understanding how to put this together!)

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

(35 / s) + [14 / (s - 5)] = 2

Answer by stanbon(75887)

During the first part of a trip, a canoeist travels 35 miles at a certain speed. The canoeist travels 14 miles on the second part or the trip at a speed 5 mph slower. The total time for the trip is 2 hours. What is the speed on each part of the trip?
----
1st leg DATA:
distance = 35 miles ; rate = x mph ; time = d/r = 25/x hrs
--------------------------
2nd leg DATA:
distance = 14 miles ; rate = x-5 mph ; time = d/r = 14/(x-5) hrs
------
Equation:
time + time = 2 hrs
25/x + 14/(x-5) = 2 hrs
25(x-5) + 14x = 2x(x-5)
25x - 125 + 14x = 2x^2 - 10x
2x^2 -49x + 125 = 0
x = 21.607 mph (rate on 1st leg)
x-5 = 16.607 mph (rate on 2nd leg)
=====================================
cheers,
Stan H.
=============